(a) Find the z-transform of (i) x[n] = au[n] +bu[n] + cºul-n – 1], lal <151 < le|| (ii) x[n] = n*au[n] (iii) x[n] = en* [c
Answer #1

Š aucnizh (a) in acno ch buny) . rucnot mum) {lal=1612160} # for a u cn) :- X (2) -n x(0.Z - un) = nzo nco = X(z) = an zin =t for cuc-n-15.- Z n(n). on Z uc-n-uz uconago = 1 - 2 uc n-> cm 2 X (2) È can ; ; replacining m-00 » X (2) em z Ž (2²2) -mazin) -n n² an uan) Differentiation property of z-transform - : 200 = R. X (2) x (no i Roc= n.am - zdxz) dz Z an ucho Z-a I-92